# Comparing the latent features of universal machine-learning interatomic potentials

### v1

@@ III.6 Local to global features
While most uMLIPs strictly work at the level of atom-centered local environments, one often needs to operate at the molecular or structural level, e.g., to build data-driven maps of datasets or to analyze the outcome of simulations trib+12pnas; isay+15cm. For these applications, it is customary to average the atomic descriptors to obtain global features Guo2022; Cheng2020, which, however, entirely disregards the variability across the atomic environments and results in significant information loss. To mitigate this and capture comprehensive structural information, we investigate the use of higher-order cumulants of atomic features to construct structure-level descriptors.
@@
where the power is applied entry-wise to each feature. The cumulants $\boldsymbol{\kappa}^{(k)}_{S}$ are then derived as follows (full formulas up to order 8 are provided in the SI XI):
@@
Figure 8: Reconstruction errors of the progressive cumulative structure-level features, constructed as concatenations of central moments of last-layer features per atom. The features are derived from PET-MAD on the MAD test. $\phi_{S}^{(m)}$ denote cumulative descriptors up to order $m$. For LFRE, we set the neighborhood size $k$ to the largest feature dimension among the compared pair features.
Applying eight-order cumulative descriptors to cross-model analysis on the same MAD test dataset (SI XI Fig. S9) yields patterns similar to atomic-level features (Fig. 1) but with amplified dissimilarities. Overall, the errors are much higher (on average, the GFRE is now 0.87, and LFRE is 0.83). This pattern suggests that the differences in the model latent features arising from the rare or asymmetric atomic-environment contributions is better captured by the higher cumulants. Practically, if the goal is to discriminate subtle model embeddings and emphasize rare motifs, inclusion of the higher cumulants is highly recommended.
To further illustrate the impact of higher-order statistics, we compare PCovR projections of structure-level descriptors built from mean-only features ($\tilde{\boldsymbol{\kappa}}_{S}^{(1)}$) with those using full eighth-order cumulative cumulants ($\phi_{S}^{(8)}$) (Fig. 9). When only the mean atomic features are used (top row), the projections exhibit a loosely consistent motif, characterized by a linear cluster corresponding to a few of the subsets and then the remaining subsets dispersed around it. However, even at this level, the exact shape and subset assignment around the projection motif vary across the different models. For the eighth-order cumulant descriptors (middle row), the linear cluster fragments disappears entirely for MACE-mp-03b, PET-MAD, and PDA-3.1, and the overall point distributions become highly model-specific with almost no cross-model alignment. Such vast differences are consistent with the higher values of the GFRE and LFRE observed between the uMLIPs in SI XI Fig. S9.
Figure 9: PCovR projections of structure-level descriptors for different uMLIPs on the MAD test set. Top row: projections using mean atomic features ($\tilde{\boldsymbol{\kappa}}_{S}^{(1)}$), colored by MAD subsets. Middle row: projections using eighth-order cumulative cumulant features ($\phi_{S}^{(8)}$), colored by MAD subsets. Bottom row: same eighth-order cumulant features colored by cohesive energy.
@@ IV Conclusions
The rapidly growing number of “universal” atomistic machine learning potentials and models, differing in architecture and training strategy but capable of achieving similar accuracies on benchmarks, raises questions on how different models achieve their learning targets, especially in terms of their internal representation of atomic structures. In this work, we have carried out systematic comparisons of the latent spaces of several universal ML models and their variants, using feature reconstruction error metrics that quantitatively compare the “information content” of different sets of features. We see that the reconstruction errors are generally high between the latent features of different uMLIPs, suggesting that there are significant differences between the internal representations of the models despite their comparable accuracy. The local feature reconstruction errors are lower, indicating the existence of nonlinear relations between the features of different models. They are nevertheless substantial, which indicates that different model architectures choose to retain different structural and chemical information.
When comparing the variants of the same model architecture trained on different “universal datasets”, the trend becomes highly dependent on the adopted strategy, with single-task and multi-head models exhibiting higher similarities than a model trained with the MoLE. We also note that the model variant corresponding to the large OMat24 dataset generally exhibits the smallest reconstruction errors relative to the latent features obtained with smaller, more specific training sets. This trend is consistent among all architectures, which indicates that a large and diverse dataset is not only beneficial for pre-training, but also encourages the learning of richer internal representations. Similar trend is observed when comparing the PET-MAD uMLIP with PET-MAD-DOS, a model built from the same architecture and training set but targeting the electron density of states, which contains energy-resolved targets and yields more information-rich features.
@@
The comparison between the backbone and last-layer features reveals that the backbone layer contains relatively more information, which affirms the multi-head strategy of creating multiple heads for different prediction targets. We also see large information asymmetry between the BB and LL features of PET-MAD, which we trace to the fact that deep layers of an unconstrained model contain a non-invariant character that is projected out as the model approaches the prediction of the scalar energy target. Lastly, we have extended the atomic latent features to structure-level descriptors via higher-order cumulants. The higher cumulants fully subsume lower-order statistics, indicating that they can preserve information that is lost when using average descriptors to represent structures at the global level, which is a common practice in the field.
Our study demonstrates that feature reconstruction errors are useful quantitative tools in comparing and contrasting the internal representations of different uMLIPs. We envision that these metrics could prove helpful in designing better models or devising new training strategies. For example, one could imagine using the GFRE and LFRE to guide the optimization of architectures and hyperparameters to maximize the descriptive power of the inner representations. The size and diversity of the training set, as well as the learning targets, also promote learning more informative latent features, which could then accelerate the training of domain-specific models via fine-tuning and transfer learning. Based on our results, we argue that accuracy alone does not define a atomistic ML model. By exposing the hidden diversity of uMLIP representations, we offer a foundation for more transparent and robust design of atomistic models.
@@
The code and data required to reproduce the results of this work are available as a record matcloud_record within the Materials Cloud Talirz_2020 Archive (DOI: 10.24435/materialscloud:gb-5z). Model weights used in this work were obtained from their public repositories: mace-foundations (https://github.com/ACEsuit/mace-foundations), PET-MAD (https://github.com/lab-cosmo/pet-mad), DPA-3.1-3M (https://www.aissquare.com/models/detail?name=DPA-3.1-3M&id=343&pageType=models), and fairchem (https://fair-chem.github.io/). Original datasets can be obtained from the respective cited sources.
@@ V Feature space comparison metrics
To aid intuition for the difference between the Global Feature Reconstruction Error (GFRE) and the Local Feature Reconstruction Error (LFRE) introduced in Sec. II.1, we present a minimal illustrative example in Fig. S1.
Figure S1: A schematic overview of the global (GFRE) and local (LFRE) feature reconstruction errors. Panel (A) shows a one-dimensional manifold of atomic environments embedded into two two-dimensional feature spaces, $F$ and $F^{\prime}$, related by a nonlinear transformation. Panel (B) illustrates $\mathrm{GFRE}(F,F^{\prime})$, where a single global linear mapping $\widehat{P}^{FF^{\prime}}$ is applied to reconstruct $F^{\prime}$ from $F$. The dashed red line indicates the globally optimal linear fit $\widehat{P}^{FF^{\prime}}X^{F}$. Panel (C) shows $\mathrm{LFRE}(F,F^{\prime})$: for each atomic environment $i$, a separate local linear mapping $\widehat{P}^{FF^{\prime}}_{(i)}$ is fitted to its $k$ nearest neighbors $\mathcal{N}_{i}$ in feature space $F$.
@@ VI Latent feature comparison between uMLIPs on Alexandria dataset
Figure S2: Heat maps of GFRE (left) and LFRE (right) for atomic last-layer latent features of MACE-MP-03b, PET-MAD, DPA-3.1, and UMA-S-1P1, computed with the atomic environments sampled from the subset of the Alexandria dataset. Each cell represents the reconstruction error when mapping latent features of the “source” (row) to “target” (column).
To ensure independence from training distributions and mitigate potential biases, we evaluate reconstruction errors on the subsampled Alexandria dataset (sAlexandria) omat24; Alexandria, which contains a broad coverage of inorganic materials across three-, two-, and one-dimensional periodic structures. We first randomly select 20,000 structures from the validation split, then filter to exclude the single-atom structures and structures with no interatomic distances below 6 Å (incompatible with UMA), and those containing actinide elements (as atomic numbers 89 – 94 are not supported by MACE), yielding 169,836 atomic environments from 16,306 structures. From these, 1000 atomic environments are randomly sampled to comprise $\mathcal{D}_{\mathrm{test}}$, and all remaining environments are used as $\mathcal{D}_{\mathrm{train}}$ to obtain the regression weights $\widehat{P}_{FF^{\prime}}$.
@@
VII Comparing model variants
VII.1 Multi-task handling
VII.1.1 DPA-3.1
DPA-3.1 incorporates a multi-task framework with a shared backbone network conditioned on dataset identifiers via one-hot encoding DPA3. We compare its branches corresponding to OMat24, MPtrj, OC20, ODAC, and SPICE omat24; MPTRAJ; OC20; ODAC; SPICE.
As shown in Fig. S3, the latent spaces of these branches are remarkably consistent despite the extreme diversity of the underlying data. LFRE remains particularly low (with the maximum of 0.14), indicating highly similar local neighborhoods across branches. This suggests that this multi-task design with full backbone sharing enforces highly similar latent representations across branches.
@@ VII.1 Multi-task handling
VII.1.2 Universal PET checkpoints
We evaluate mutual reconstruction between PET universal models trained on different datasets. Table S1 summarizes the characteristics of the five analyzed PET variants, including model size and training protocol.
Table S1: Overview of PET model variants used in the reconstruction error analysis. All “O”-prefixed models are pre-trained on OMat24 before training on the target dataset. sAlex stands for the subsampled Alexandria dataset (sAlexandria) omat24; Alexandria.
@@ VII.1.2 Universal PET checkpoints
Figure S4: Reconstruction errors among PET variants trained on MAD, OAM, OMAT, and OMATPES on the MAD test omat24; MATPES; MPTRAJ; MAD.
From Fig. S4, MAD and OAM checkpoints are the most challenging to reconstruct with global linear approximation. OMat24 pretrained models (OMAD, OMAT, and OMATPES) show slightly lower reconstruction, sharing the same latent space foundation from the pre-training.
VII.2 Dataset domain
We further probe the effect of training data by comparing medium-sized models, MACE-MP-03b MACE, trained on MPtrj MPTRAJ, a materials-focused dataset, and MACE-OFF23 MACEOFF, trained on SPICE SPICE, an organic systems-focused dataset.
Reconstruction errors are evaluated on the organic subset of the MAD test set. From Table S2, the materials-trained model (MP-03b) can reconstruct the organic-trained model almost perfectly (GFRE = 0.02). In the reverse direction, the organic-trained model fails to reconstruct the materials model globally (GFRE = 0.36), despite near-perfect latent-feature matching (LFRE = 0.04). This suggests that the latent feature spaces learned by the two models overlap extremely well on organic geometries. The global reconstruction from OFF23 to MP-03b diverges, as the materials model was trained on a vastly more diverse set, and has learned a richer global mapping that the organic-only model cannot replicate when forced to use the materials model’s latent representations.
@@ VII.2 Dataset domain
VII.3 Model sizes
VII.3.1 MACE-MP-0 and MACE-OFF23
We examine how well MACE models of different sizes (small, medium, large) can reconstruct the last-layer features of one another on the organic structures of the MAD dataset. In Fig. S5, the upper triangle shows results for the MPtrj-trained MACE-MP-0a evaluated on the MAD test set, the lower triangle for the organic-molecule family model, MACE-OFF23 MACEOFF, evaluated on the organic subset of MAD test.
@@ VII.3.1 MACE-MP-0 and MACE-OFF23
Notably, MACE-MP-0a checkpoints reconstruct each other almost perfectly, regardless of the model size. For MACE-OFF23, the difference appears when a small or medium model tries to reconstruct a large one (GFRE is 0.48, LFRE is 0.31 on average between them), meaning the larger model contains expressive capacity that smaller models cannot fully replicate. It could stem from the differences between large variants in hyperparameters or training protocol.
VII.3.2 PET-OMAT
We compare the PET-OMAT checkpoints Bigi2025prep of different sizes, as they can be used as a pre-training base for fine-tuning. The model sizes depend on the number of GNN layers ($N_{\text{gnn}}$), the number of transformer layers ($N_{\text{tl}}$), the neural network width ($d_{\text{pet}}$), and the cutoff radius ($r_{\text{cut}}$) (Table S3).
Table S3: PET-OMAT model sizes.

### v2

@@ III.6 Local to global features
While most uMLIPs strictly work at the level of atom-centered local environments, one often needs to operate at the molecular or structural level, e.g., to build data-driven maps of datasets or to analyze the outcome of simulations Tribello et al. (2012); Isayev et al. (2015). For these applications, it is customary to average the atomic descriptors to obtain global features Guo et al. (2022); Cheng et al. (2020), which, however, entirely disregards the variability across the atomic environments and results in significant information loss. To mitigate this and capture comprehensive structural information, we investigate the use of higher-order cumulants of atomic features to construct structure-level descriptors.
@@
where the power is applied entry-wise to each feature. The cumulants $\boldsymbol{\kappa}^{(k)}_{S}$ are then derived as follows (full formulae up to order 8 are provided in the SI XII):
@@
Figure 8: The reconstruction errors of the progressive cumulative structure-level features, constructed as concatenations of central moments of last-layer features per atom. The features are derived from PET-MAD for the MAD test subset. $\phi_{S}^{(m)}$ denote cumulative descriptors up to order $m$. For the local feature reconstruction error (LFRE), we set the neighborhood size $k$ to the larger feature dimension between the compared feature pairs.
Applying eighth-order cumulative descriptors to cross-model analysis on the same MAD test dataset (SI XII Fig. S10) yields patterns similar to atomic-level features (Fig. 1) but with amplified dissimilarities. Overall, the errors are much higher (on average, the GFRE is now 0.87, and LFRE is 0.83). This pattern suggests that the differences in the model latent features arising from the rare or asymmetric atomic-environment contributions is better captured by the higher cumulants. Practically, if the goal is to discriminate subtle model embeddings and emphasize rare motifs, inclusion of the higher cumulants is highly recommended.
To further illustrate the impact of higher-order statistics, we compare PCovR projections of structure-level descriptors built from mean-only features ($\tilde{\boldsymbol{\kappa}}_{S}^{(1)}$) with those using full eighth-order cumulative cumulants ($\phi_{S}^{(8)}$) (Fig. 9). When only the mean atomic features are used (top row), the projections exhibit a loosely consistent motif, characterized by a linear cluster corresponding to a few of the subsets and then the remaining subsets dispersed around it. Even at this level, however, the exact shape and subset assignment around the projection motif vary across the different models. For the eighth-order cumulant descriptors (middle row), the linear cluster fragments disappear entirely for MACE-MP-0b3, PET-MAD, and DPA-3.1, and the overall point distributions become highly model-specific with almost no cross-model alignment. Such vast differences are consistent with the higher values of the GFRE and LFRE observed between the uMLIPs in SI XII Fig. S10.
Figure 9: The PCovR projections of structure-level descriptors for different uMLIPs on the MAD test subset. Top row: projections using mean atomic features ($\tilde{\boldsymbol{\kappa}}_{S}^{(1)}$), colored by the different structural subsets of the MAD dataset. Middle row: projections using eighth-order cumulative cumulant features ($\phi_{S}^{(8)}$), again colored by respective structural subsets. Bottom row: same eighth-order cumulant features colored by the structural cohesive energy used as target in PCovR analysis.
@@ IV Conclusions
The rapidly growing number of “universal” atomistic machine learning potentials and models, differing in architecture and training strategy but capable of achieving similar accuracies on benchmarks, raises questions of how the different models achieve their learning targets, especially in terms of their internal representations of the chemical systems. In this work, we have carried out systematic comparisons of the latent spaces of several universal ML models and their variants, using the feature reconstruction error metrics that quantitatively compare the “information content” of different sets of features. We see that the reconstruction errors are generally high between the latent features of different uMLIPs, revealing the presence of significant differences between the internal representations of the models despite their comparable accuracy. The local feature reconstruction errors are lower, which indicate the existence of nonlinear relations between the features of different models. They are nevertheless substantial, which suggests that different models choose to retain distinct structural and chemical information.
When comparing the variants of the same model architecture trained on different “universal” datasets, the trend becomes highly dependent on the adopted training strategy, with single-task and multi-head models exhibiting higher similarities than a model trained with the MoLE. We also note that the model variant corresponding to the large OMat24 dataset generally exhibits the smallest reconstruction errors compared to the other cases. This hints that a large, diverse dataset is not only beneficial for pre-training, but also encourages the learning of richer internal representations. A similar trend is observed when comparing the PET-MAD with PET-MAD-DOS, a model built from the same architecture and training set but targeting the energy-resolved electronic density of states, thereby yielding more information-rich features.
@@
The comparison between the backbone and last-layer features reveals that the backbone layer contains relatively richer information, which retrospectively explains the success of using multiple heads for different prediction targets. In the particular case of PET-MAD, a rotationally unconstrained model, we further attribute the large information asymmetry to the fact that the deeper layers of PET would exhibit a non-invariant character that gets successively projected out as the model approaches the prediction of the scalar energy target. Lastly, we have extended the raw, atomic, local latent features of the uMLIPs to structural, global descriptors via higher-order cumulants. The higher cumulants fully subsume the lower-order statistics, which critically indicate that significant information becomes lost when using the averages of atomic features as structural descriptors, which is a common practice in the field.
Our results yield several practical insights for atomistic ML practitioners. In particular, one can apply the feature reconstruction error metrics to understand the information flow within a model architecture, for example, by quantifying how much generality is lost in the readout layers. One could also imagine using the GFRE and LFRE to guide the optimization of architectural hyperparameters to maximize the descriptive power of internal representations. In the context of fine-tuning, the changes in the latent features measured via the FREs can be linked to catastrophic forgetting and used to monitor the adjustments of the internal representations and prevent any significant loss of generalizability to out-of-distribution configurations. Given the hidden diversity of the uMLIP latent spaces uncovered in our study, we argue that predictive accuracy alone cannot sufficiently characterize atomistic ML models. Through the adoption of feature reconstruction error metrics, we establish a principled foundation for a more transparent, interpretable, and robust design of future atomistic ML models.
@@
The code and data required to reproduce the results of this work are available as a record Chorna et al. (2025) within the Materials Cloud Talirz et al. (2020) Archive (DOI: 10.24435/materialscloud:r5-vh). Model weights used in this work were obtained from their public repositories: mace-foundations (https://github.com/ACEsuit/mace-foundations), PET-MAD (https://github.com/lab-cosmo/pet-mad), DPA-3.1-3M (https://www.aissquare.com/models/detail?name=DPA-3.1-3M&id=343&pageType=models), and fairchem (https://fair-chem.github.io/). Original datasets can be obtained from the respective cited sources.
@@ V Feature space comparison metrics
To aid the reader’s understanding of the the difference between the global feature reconstruction error (GFRE) and the local feature reconstruction error (LFRE) introduced in Sec. II.1, we present a minimal illustrative example in Fig. S1.
Figure S1: A schematic overview of the global (GFRE) and local (LFRE) feature reconstruction errors. (A) shows a one-dimensional manifold of atomic environments embedded into two two-dimensional feature spaces, $F$ and $F^{\prime}$, related by a nonlinear transformation. (B) illustrates $\mathrm{GFRE}(F,F^{\prime})$, where a single global linear mapping $\widehat{P}^{FF^{\prime}}$ is applied to reconstruct $F^{\prime}$ from $F$. The dashed red line indicates the globally optimal linear fit $\widehat{P}^{FF^{\prime}}X^{F}$. (C) shows $\mathrm{LFRE}(F,F^{\prime})$: for each atomic environment $i$, a separate local linear mapping $\widehat{P}^{FF^{\prime}}_{(i)}$ is fitted to its $k$ nearest neighbors $\mathcal{N}_{i}$ in feature space $F$.
@@ VI Latent feature comparison between uMLIPs on Alexandria dataset
Figure S2: The reconstruction errors for atomic last-layer latent features of MACE-MP-03b, PET-MAD, DPA-3.1, and UMA-S-1P1, computed for the atomic environments sampled from the sAlex dataset.
To ensure independence from training distributions and mitigate potential biases, we evaluate reconstruction errors on the subsampled Alexandria dataset (sAlex) Barroso-Luque et al. (2024); Schmidt et al. (2024), which contains a broad coverage of inorganic materials across three-, two-, and one-dimensional periodic structures. We first randomly select 20,000 structures from the validation split, then filter to exclude the single-atom structures and structures with no interatomic distances below 6 Å (incompatible with UMA-S-1P1), and those containing actinide elements (as atomic numbers 89–94 are not supported by MACE-MP-0b3), yielding 169,836 atomic environments from 16,306 structures. From these, 1000 atomic environments are randomly sampled to comprise $\mathcal{D}_{\mathrm{test}}$, and all remaining environments are used as $\mathcal{D}_{\mathrm{train}}$ to obtain the regression weights $\widehat{P}_{FF^{\prime}}$.
@@
VII Comparing backbone features of uMLIPs
The lower reconstruction errors for backbone features (average off-diagonal GFRE 0.54 vs 0.66 for last-layer, LFRE 0.30 vs 0.37) indicate that representations immediately after message-passing are more consistent and linearly reconstructable.
Figure S3: The reconstruction errors for atomic backbone features computed across MACE-MP-0b3, PET-MAD, DPA-3.1, and UMA-S-1P1, evaluated on the MAD test subset. Compared to last-layer features (Fig. 1), backbone representations exhibit consistently lower cross-model errors.
VIII Comparing model variants
VIII.1 Multi-task handling
VIII.1.1 DPA-3.1
DPA-3.1 incorporates a multi-task framework with a shared backbone network conditioned on dataset identifiers via one-hot encoding Zhang et al. (2025). We compare its branches corresponding to OMat24, MPtrj, OC20, ODAC, and SPICE Barroso-Luque et al. (2024); Deng et al. (2023); Chanussot et al. (2021); Sriram et al. (2023); Eastman et al. (2023). As shown in Fig. S4, the latent spaces of these branches are remarkably consistent despite the extreme diversity of the underlying data. LFRE remains particularly low (with the maximum of 0.14), indicating highly similar local neighborhoods across branches. This suggests that this multi-task design with full backbone sharing enforces highly similar latent representations across branches.
@@ VIII.1 Multi-task handling
VIII.1.2 Universal PET checkpoints
Here, we evaluate the feature reconstruction errors across single-task PET universal models trained on different datasets. Table S1 summarizes the characteristics of the five PET variants analyzed, including model size and training protocol. All models were obtained from the public UPET repository Laboratory of Computational Science and Modeling (COSMO), EPFL (2026). From Fig. S5, MAD and OAM checkpoints are the most challenging to reconstruct with global linear approximation. OMat24 pretrained models (OMAD, OMAT, and OMATPES) show slightly lower reconstruction, sharing the same latent space foundation from the pre-training.
Table S1: Overview of PET model variants used in the reconstruction error analysis. All “O”-prefixed models are pre-trained on OMat24 before training on the target dataset. sAlex stands for the subsampled Alexandria dataset (sAlexandria) Barroso-Luque et al. (2024); Schmidt et al. (2024).
@@ VIII.1.2 Universal PET checkpoints
Figure S5: Reconstruction errors among PET variants trained on MAD, OAM, OMAT, and OMATPES on the MAD test Barroso-Luque et al. (2024); Kaplan et al. (2025); Deng et al. (2023); Mazitov et al. (2025b).
VIII.2 Dataset domain
We further probe the effect of training dataset domain by comparing medium-sized models, MACE-MP-0b3 Batatia et al. (2025a) trained on MPtrj Deng et al. (2023), a materials-focused dataset, and MACE-OFF23 Kovács et al. (2025) trained on SPICE Eastman et al. (2023), an organics-focused dataset.
Reconstruction errors have been computed for only the organic structures found within the MAD test set. From Table S2, MACE-MP-0b3 can reconstruct the organic-trained model almost perfectly (GFRE = 0.02). In the reverse direction, the MACE-OFF23 fails to reconstruct the MACE-MP-0b3 features globally (GFRE = 0.36), despite near-perfect local feature reconstruction (LFRE = 0.04). This is explained by the fact that MACE-MP-0b3 was trained on a vastly more diverse dataset, and has learned a richer global mapping that MACE-OFF23 simply cannot replicate due to a significantly smaller dataset scope.
@@ VIII.2 Dataset domain
VIII.3 Model sizes
VIII.3.1 MACE-MP-0a and MACE-OFF23
In this section, we examine how well the MACE models of different sizes (small, medium, large) can reconstruct the last-layer features of one another. In Fig. S6, the upper triangle shows results for the MPtrj-trained MACE-MP-0a evaluated on the MAD test subset, and the lower triangle shows the results for MACE-OFF23 Kovács et al. (2025) evaluated on the organic structures found in the MAD test subset.
@@ VIII.3.1 MACE-MP-0a and MACE-OFF23
Notably, MACE-MP-0a checkpoints reconstruct each other almost perfectly, regardless of the model size. For MACE-OFF23, error rises when a small or medium model tries to reconstruct the large one (GFRE is 0.48, LFRE is 0.31 on average between them), meaning the larger model contains expressive capacity that smaller models cannot fully replicate. These distinct trends could stem from the differences in the hyperparameters or training protocol between MACE-MP series and MACE-OFF, and/or the differences in the scope of the training datasets.
VIII.3.2 PET-OMAT
In addition to MACE variants, we also compare the PET-OMAT checkpoints Laboratory of Computational Science and Modeling (COSMO), EPFL (2026) of different sizes. The model sizes depend on the combination of the number of GNN layers ($N_{\text{gnn}}$), the number of transformer layers ($N_{\text{tl}}$), the neural network width ($d_{\text{pet}}$), and the cutoff radius ($r_{\text{cut}}$) (see Table S3).
Table S3: PET-OMAT model hyperparameters.